# Does DetectGPT Fully Utilize Perturbation? Selective Perturbation on Model-Based Contrastive Learning Detector would be Better

### 01 Feb 2024 v1

@@ Abstract
The burgeoning capabilities of large language models (LLMs) have raised growing concerns about abuse. DetectGPT Mitchell et al. (2023), a zero-shot metric-based unsupervised machine-generated text detector, first introduces perturbation and shows great performance improvement. However, DetectGPT’s random perturbation strategy might introduce noise, limiting the distinguishability and further performance improvements. Moreover, its logit regression module relies on setting the threshold, which harms the generalizability and applicability of individual or small-batch inputs. Hence, we propose a novel detector, Pecola, which uses selective strategy perturbation to relieve the important information loss caused by random masking, and multi-pair contrastive learning to capture the implicit pattern information during perturbation, facilitating few-shot performance. The experiments show that Pecola outperforms the SOTA method by 1.20% in accuracy on average on four public datasets. We further analyze the effectiveness, robustness, and generalization of our perturbation method.
@@ 1 Introduction
Machine-generated text (MGT) detection is to discriminate MGT from human-written texts (HWT), preventing abuse of large language models (LLMs), including academic misconduct Vasilatos et al. (2023), spam synthesis Dou et al. (2020), untrustworthy news Zellers et al. (2019), etc. Currently, existing MGT detection methods can be mainly classified into two genres Wu et al. (2023), i.e., training model-based Liu et al. (2023); Hu et al. (2023); Verma et al. (2023); OpenAI (2023) and zero-shot metric-based methods Gehrmann et al. (2019); Mitchell et al. (2023); Yang et al. (2023); Bao et al. (2023). In general terms, training model-based methods can achieve better accuracy than zero-shot metric-based methods, especially generalizable to black-box generators, but are more costly during data collection, fine-tuning, and running. Moreover, zero-shot metric-based methods show better interpretability than training model-based ones.
@@
However, DetectGPT still has three significant defects: (i) DetectGPT is fragile to be generalized on zero-shot settings due to its high dependency on the threshold of the logit regression module. Further, its application scenario is thus limited to large batch inputs but fails on individual inputs. (ii) DetectGPT does not fully utilize the perturbation. As a metrics-based method, it only considers the probability difference caused by perturbation, which is overly simplified and slightly indistinguishable. Perturbation should indeed be a stronger augment that carries implicit language pattern information. (iii) DetectGPT perturbs the original sample randomly and unrestricted, which could introduce more noise and negatively impact the performance Kim et al. (2022). For example, Liu et al. (2023) find entity-relationship plays a role in the detection, which might be destroyed in random perturbation of DetectGPT.
In this paper, we thus propose a Perturbation-based Contrastive Learning model, Pecola, for MGT detection, toward the defects with two stages, i.e., Selective Strategy Perturbation and Token-Level Weighted Multi-Pairwise Contrastive Learning. Firstly, Selective Strategy Perturbation is a token-level rewriting method with restrictions on modifying important texts Campos et al. (2020) to reduce noise. The motivation is to simulate human modification Verma and Lee (2017); Fetaya et al. (2020); Wang et al. (2019). The perturbation strategy consists of removing and substituting tokens, as shown in Fig. 1. The experiments show that our Selective Strategy Perturbation method can improve the performance of both metrics-based (i.e., DetectGPT) and model-based methods. Secondly, we use a Multi-Pairwise Contrastive Learning model to process the perturbed texts. Different from the logit regression module in DetectGPT, the trained model is generalizable without any threshold setting, and it can deal with individual inputs. Moreover, by utilizing multi-pairwise contrastive learning, the model could better utilize perturbation to focus on the language pattern gap between HWT and MGT. The importance weight from the perturbation stage is also reused as contrastive learning weight. Notably, by using contrastive learning, Pecola is a strong few-shot training model-based method, which is more data-efficient during training than existing training-based methods. Finally, extensive experiments show Pecola is significantly superior to baseline and SOTA methods on four datasets, e.g., Pecola surpasses SOTA by 1.20% in accuracy on average under few-shot settings. Further experiments show that Pecola is better on generalization, robustness, and effectiveness.
@@
Selective Perturbation: We propose a novel Selective Strategy Perturbation method considering token importance, which reduces the noise of random in DetectGPT and enhances its benefits to both supervised and unsupervised approaches.
Contrastive Learning: We propose a contrastive learning model to replace the logit regression of DetectGPT, which frees the detector from setting the threshold, enables it to deal with individual input, and could be generalizable and effective on the few-shot setting by contrast perturbed texts with origin ones.
Outperformance: Our detector Pecola outperforms all compared models on four public datasets. And Pecola is more robust to the choice of base model and filling model.
@@ 2 Related Work
Machine-generated Text Detection. While fine-tuning pre-trained models have proven effective for MGT detection Wahle et al. (2022); Hu et al. (2023), the requirement for annotated datasets poses a significant challenge due to the proliferation of unchecked, high-quality generated texts. To address this challenge, DetectGPT Mitchell et al. (2023) has demonstrated strong performance in white-box zero-shot settings. Similarly, CoCo Liu et al. (2023) is designed to detect MGT with low resource annotations, utilizing a coherence-based contrastive learning model. Recently, the Watermark method Kirchenbauer et al. (2023) has been introduced to mitigate the risk associated with unchecked MGTs by embedding imperceptible signals within text outputs. In contrast to previous methods, our approach integrates data perturbation with contrastive learning, with a particular emphasis on reducing reliance on mask-filling models and enhancing performance in few-shot scenarios.
Perturbation. Data perturbation methods find frequent application in text classification tasks Gao et al. (2022); Shum et al. (2023), which is commonly employed through the technique of consistency regularization Xie et al. (2020); Chen et al. (2020). Nevertheless, in MGT detection, previous perturbation methods have exhibited certain limitations. For instance, they often resort to randomly selecting target tokens for synonym replacement Wang et al. (2018), deletion, or insertion Wei and Zou (2019), and fine-tuning pre-trained models to fill text spans of variable lengths Gao et al. (2022). While these methods do enhance text diversity, the indiscriminate replacement of tokens without guided rules can lead to the generation of less reliable texts. These limitations motivate us to devise data perturbation methods tailored for MGT detection. Our approach, guided selectively, aims to better represent meaningful recombination spaces while preserving the inherent semantic features of the text, ultimately enhancing the diversity of our samples.
@@ 3 Methodology
The workflow of Pecola mainly consists of two stages: Selective Strategy Perturbation and Supervised Contrastive Learning for fine-tuning PLM, as shown in Fig. 2.
@@ 3.1 Selective Strategy Perturbation
Token Mask Selection. After getting the important tokens set of each text input , we use special token [MASK] to replace some of the tokens in the text input to construct masked perturbation input . In order to relieve the information loss caused by masking perturbation, we add regularization to the traditional random masking method and use a selective masking strategy to prevent important tokens from being masked.
@@ 3.2 Token-Level Weighted Multi-Pairwise Contrastive Learning
Importance-based Feature Reconstruction. Existing MGT methods Liu et al. (2023) often uniformly extract all token information in the text, ignoring the huge impact of a few important tokens on the detection model. In this work, we reconstruct the token feature extracted by PLM according to the importance of the token in the input text, allowing the detection model to focus more on important token information.
We assign adaptive weights to all tokens in the input:
@@
Multi-Pairwise Contrastive Learning. Considering that existing works Gunel et al. (2021); Zhou et al. (2021); Liu et al. (2023) mainly concentrate on single-input feature learning while overlooking input correlations, we introduce contrastive learning into MGT. This approach enables our model to discern the distinct featurinputes of variously labeled data, more accurately capture input features, and significantly enhance performance in few-shot setting.
@@
| Dataset | Metric | Shot | GPT-2 | IT:Clust | RoBERTa | CE+SCL | CE+Margin | CoCo | GLTR | DetectGPT | Pecola |
| Grover | Acc | 32 | 47.71 | 41.57 | 48.83 | 55.86 | 56.79 | 51.60 | 56.61 | 55.02 | 59.03 |
| 64 | 49.45 | 46.45 | 56.88 | 57.57 | 58.92 | 58.27 | 56.61 | 54.61 | 60.94 | ||
| 128 | 51.67 | 50.72 | 59.28 | 60.33 | 60.44 | 58.97 | 58.48 | 55.78 | 63.60 | ||
| 512 | 53.24 | 56.08 | 70.39 | 72.38 | 72.15 | 70.07 | 62.26 | 55.56 | 73.12 | ||
| F1 | 32 | 41.56 | 40.79 | 44.13 | 51.56 | 53.21 | 47.33 | 52.77 | 51.09 | 53.95 | |
| 64 | 46.48 | 46.10 | 52.88 | 53.39 | 54.99 | 44.70 | 52.77 | 48.07 | 55.48 | ||
| 128 | 46.70 | 51.37 | 54.69 | 55.74 | 55.54 | 51.44 | 54.47 | 53.78 | 58.98 | ||
| 512 | 47.58 | 51.80 | 64.49 | 67.02 | 66.25 | 65.15 | 57.11 | 53.32 | 68.24 | ||
| GPT-2 | Acc | 32 | 51.20 | 51.02 | 70.53 | 69.32 | 70.00 | 71.69 | 75.99 | 68.59 | 75.42 |
| 64 | 51.89 | 54.32 | 74.41 | 73.77 | 74.04 | 73.20 | 75.76 | 71.12 | 78.92 | ||
| 128 | 54.68 | 59.66 | 79.77 | 80.18 | 80.93 | 79.44 | 75.77 | 71.74 | 82.58 | ||
| 512 | 66.99 | 71.59 | 84.07 | 84.76 | 84.89 | 84.30 | 75.86 | 71.74 | 85.75 | ||
| F1 | 32 | 43.22 | 48.45 | 66.57 | 64.89 | 69.89 | 71.19 | 72.45 | 65.50 | 75.10 | |
| 64 | 47.14 | 53.87 | 73.91 | 72.32 | 73.94 | 69.79 | 70.87 | 66.58 | 78.88 | ||
| 128 | 53.09 | 59.48 | 79.49 | 80.00 | 80.79 | 76.10 | 71.16 | 66.13 | 82.54 | ||
| 512 | 66.99 | 70.42 | 84.01 | 84.72 | 84.86 | 83.88 | 75.56 | 66.13 | 85.72 | ||
| GPT-3.5 | Acc | 32 | 68.79 | 61.82 | 90.54 | 92.44 | 92.85 | 93.27 | 92.55 | 84.42 | 95.80 |
| 64 | 82.58 | 77.70 | 96.85 | 96.86 | 97.32 | 95.76 | 91.00 | 82.58 | 98.01 | ||
| 128 | 86.20 | 92.54 | 97.50 | 98.00 | 98.00 | 96.26 | 91.60 | 85.33 | 98.06 | ||
| 512 | 92.81 | 98.13 | 98.97 | 98.99 | 98.92 | 98.05 | 92.60 | 85.57 | 99.14 | ||
| F1 | 32 | 66.58 | 60.95 | 90.27 | 92.42 | 92.81 | 92.72 | 92.71 | 84.43 | 95.80 | |
| 64 | 82.47 | 77.33 | 96.84 | 96.86 | 97.47 | 95.45 | 91.49 | 86.16 | 98.01 | ||
| 128 | 86.17 | 92.50 | 97.50 | 98.00 | 98.00 | 97.57 | 91.96 | 86.13 | 98.06 | ||
| 512 | 92.79 | 98.13 | 98.97 | 98.99 | 98.92 | 97.88 | 92.71 | 86.20 | 99.14 | ||
| HC3 | Acc | 32 | 61.80 | 77.00 | 93.36 | 95.33 | 95.46 | 92.11 | 97.30 | 94.54 | 97.19 |
| 64 | 82.51 | 91.69 | 96.97 | 97.81 | 97.81 | 95.50 | 98.13 | 95.03 | 98.59 | ||
| 128 | 93.54 | 95.43 | 97.56 | 98.17 | 98.14 | 97.57 | 98.29 | 95.10 | 98.63 | ||
| 512 | 96.62 | 97.98 | 98.85 | 98.93 | 98.99 | 98.58 | 98.31 | 95.13 | 99.15 | ||
| F1 | 32 | 52.47 | 76.47 | 93.34 | 95.32 | 95.45 | 92.07 | 97.30 | 94.29 | 97.19 | |
| 64 | 79.57 | 91.67 | 96.97 | 97.81 | 97.81 | 95.50 | 98.12 | 94.95 | 98.59 | ||
| 128 | 93.51 | 95.43 | 97.56 | 98.17 | 98.14 | 97.59 | 98.29 | 95.01 | 98.63 | ||
| 512 | 96.62 | 97.98 | 98.85 | 98.93 | 98.99 | 98.59 | 98.31 | 95.05 | 99.15 |

### 24 Feb 2024 v3

# Does DetectGPT Fully Utilize Perturbation? Bridge Selective Perturbation to Fine-tuned Contrastive Learning Detector would be Better

@@ Abstract
The burgeoning generative capabilities of large language models (LLMs) have raised growing concerns about abuse, demanding automatic machine-generated text detectors. DetectGPT Mitchell et al. (2023), a zero-shot metric-based detector, first introduces perturbation and shows great performance improvement. However, in DetectGPT, random perturbation strategy could introduce noise, and logit regression depends on threshold, harming the generalizability and applicability of individual or small-batch inputs. Hence, we propose a novel fine-tuned detector, Pecola, bridging metric-based and fine-tuned detectors by contrastive learning on selective perturbation. Selective strategy retains important tokens during perturbation and weights for multi-pair contrastive learning. The experiments show that Pecola outperforms the state-of-the-art by 1.20% in accuracy on average on four public datasets. And we further analyze the effectiveness, robustness, and generalization of the method.
@@ 1 Introduction
Machine-generated text (MGT) detection is to discriminate MGT from human-written texts (HWT), preventing abuse of large language models (LLMs), including academic misconduct Vasilatos et al. (2023), spam synthesis Dou et al. (2020), untrustworthy news Zellers et al. (2019), etc. Currently, existing MGT detection methods can be mainly classified into two genres Wu et al. (2023), i.e., fine-tuned methods Liu et al. (2023); Hu et al. (2023); Verma et al. (2023); OpenAI (2023) and zero-shot metric-based methods Gehrmann et al. (2019); Mitchell et al. (2023); Yang et al. (2023); Bao et al. (2024). In general terms, fine-tuned detector methods can achieve better accuracy than zero-shot metric-based methods, especially generalizable to black-box generators, but are more costly during data collection, fine-tuning, and running. On the other hand, zero-shot metric-based methods show better interpretability than fine-tuned ones.
@@
However, DetectGPT still has three significant defects: (i) DetectGPT’s reliance on the logit regression module’s threshold compromises its generalization in zero-shot settings and limited to large batch input, fails on individual inputs. (ii) DetectGPT does not fully utilize the perturbation. As a metrics-based method, it only considers the probability difference caused by perturbation, which is overly simplified and slightly indistinguishable. Perturbation should indeed be a stronger augment that carries implicit language pattern information. (iii) DetectGPT perturbs the original sample randomly and unrestricted, which could introduce more noise and negatively impact the performance Kim et al. (2022). For example, Liu et al. (2023) find entity-relationship plays a role in the detection, which might be destroyed in random perturbation of DetectGPT.
In this paper, we thus propose a Perturbation-based Contrastive Learning model, Pecola, for MGT detection, toward the defects with two stages, i.e., Selective Strategy Perturbation (Sec. 3.1) and Token-Level Weighted Multi-Pairwise Contrastive Learning (Sec. 3.2). Firstly, Selective Strategy Perturbation is a token-level rewriting method with restrictions on modifying important texts Campos et al. (2020) to reduce noise. The motivation is to simulate the human behavior of modification Verma and Lee (2017); Fetaya et al. (2020); Wang et al. (2019). The perturbation strategy consists of tokens removing and substitution, as shown in Fig. 1. The experiments show that our Selective Strategy Perturbation method can improve the performance of both metrics-based (i.e., DetectGPT) and model-based methods. Secondly, we use a Multi-Pairwise Contrastive Learning model to process the perturbed texts. Different from the logit regression module in DetectGPT, the trained model is generalizable without any threshold setting, and it can deal with individual inputs. Moreover, by utilizing multi-pairwise contrastive learning, the model could better utilize perturbation to focus on the language pattern gap between HWT and MGT. The importance weight from the perturbation stage is also reused as contrastive learning weight. Notably, by using contrastive learning, Pecola is a strong few-shot fine-tuning method, which effectively bridges and integrates metric-based and fine-tuned detector genres. Finally, extensive experiments show Pecola is significantly superior to baseline and SOTA methods on four datasets, Pecola improves by 1.20% to SOTA on average under few-shot settings, surpassing the latest methods by 3.84% among metric-based detectors and by 1.62% among fine-tuned detectors. Further experiments show that Pecola is as well better at generalization, robustness, and effectiveness.
@@
Selective Perturbation: Based on our analysis of various selective perturbation strategies, we propose a novel method considering token importance, which reduces the noise and benefits to both supervised and unsupervised approaches.
Bridge Metric and Model-based Detectors: We utilize a novel fine-tuned contrastive learning module to replace the logit regression of DetectGPT (metric-based), which frees the detector from setting the threshold, enables it to deal with individual input, and can be generalizable and effective on the few-shot setting by contrasting perturbed texts with origin ones.
Outperformance: Our detector Pecola outperforms all eight compared models on four public datasets. And Pecola is more robust to the choice of base model and filling model. Furthermore, we prove its generalization ability across domains and genres of data.
@@ 2 Related Work
Machine-generated Text Detection. While fine-tuned detectors have proven effective for MGT detection Wahle et al. (2022); Hu et al. (2023), the requirement for annotated datasets poses a significant challenge due to the proliferation of unchecked, high-quality generated texts. To address this challenge, DetectGPT Mitchell et al. (2023) and Fast-DetectGPT Bao et al. (2024) have demonstrated strong performance in white-box zero-shot settings. Similarly, CoCo Liu et al. (2023) is designed to detect MGT with low resource annotations, utilizing a coherence-based contrastive learning model. Recently, the Watermark method Kirchenbauer et al. (2023) has been introduced to mitigate the risk associated with unchecked MGTs by embedding imperceptible signals within text outputs. In contrast to previous methods, our approach integrates data perturbation with contrastive learning, with a particular emphasis on reducing reliance on mask-filling models and enhancing performance in few-shot scenarios.
Perturbation. Data perturbation methods find frequent application in text classification tasks Gao et al. (2022); Shum et al. (2023), which is commonly employed through the technique of consistency regularization Xie et al. (2020); Chen et al. (2020). Nevertheless, in MGT detection, previous perturbation methods have exhibited certain limitations. For instance, they often resort to randomly selecting target tokens for synonym replacement Wang et al. (2018), deletion, or insertion Wei and Zou (2019), and fine-tuning pre-trained models to fill text spans of variable lengths Gao et al. (2022). While these methods do enhance text diversity, the indiscriminate replacement of tokens without guided rules can lead to the generation of less reliable texts. These limitations motivate us to devise data perturbation methods tailored for MGT detection. Our approach, guided selectively, aims to better represent meaningful recombination spaces while preserving the inherent semantic features of the text, ultimately enhancing the diversity of samples.
@@ 3 Methodology
As shown in Fig. 2, the workflow of Pecola mainly consists of two stages: Selective Strategy Perturbation and Supervised Contrastive Learning for fine-tuning PLM, which enjoy the advantage of metric-based and model-based detectors, respectively.
@@ 3.1 Selective Strategy Perturbation
Mask Position Selection. After getting the important tokens set of each text input , we use special token [MASK] to replace some of the tokens in the text input to construct masked perturbation input . In order to relieve the information loss caused by masking perturbation, we add regularization to the traditional random masking method and use a selective masking strategy to prevent important tokens from being masked.
@@ 3.2 Token-Level Weighted Multi-Pairwise Contrastive Learning
Importance-based Feature Reconstruction. Existing MGT methods Liu et al. (2023) often uniformly extract all token information in the text, ignoring the huge impact of a few important tokens on the detection model. In this work, we reconstruct the token feature extracted by PLM according to the importance of the token in the input text, allowing the detection model to focus more on important token information. We assign adaptive weights to all tokens in the input:
@@
Multi-Pairwise Contrastive Learning. Considering that existing works Gunel et al. (2021); Zhou et al. (2021); Liu et al. (2023) mainly concentrate on single-input feature learning while overlooking input correlations, we introduce contrastive learning into MGT. It enables Pecola to discern the distinct featurinputes of variously labeled data, more accurately capture input features, and significantly enhance performance in few-shot setting.
@@
| Dataset | Metric | Shot | RoBERTa | GLTR | CE+SCL | CE+Margin | IT:Clust | CoCo | DetectGPT | Fast-Detect. | Pecola |
| Grover | Acc | 32 | 48.83 | 56.61 | 55.86 | 56.79 | 41.57 | 51.60 | 55.02 | 56.06 | 59.03 |
| 64 | 56.88 | 56.61 | 57.57 | 58.92 | 46.45 | 58.27 | 54.61 | 60.33 | 60.94 | ||
| 128 | 59.28 | 58.48 | 60.33 | 60.44 | 50.72 | 58.97 | 55.78 | 60.33 | 63.60 | ||
| 512 | 70.39 | 62.26 | 72.38 | 72.15 | 56.08 | 70.07 | 55.56 | 62.50 | 73.12 | ||
| F1 | 32 | 44.13 | 52.77 | 51.56 | 53.21 | 40.79 | 47.33 | 51.09 | 56.67 | 53.95 | |
| 64 | 52.88 | 52.77 | 53.39 | 54.99 | 46.10 | 44.70 | 48.07 | 57.92 | 55.48 | ||
| 128 | 54.69 | 54.47 | 55.74 | 55.54 | 51.37 | 51.44 | 53.78 | 54.89 | 58.98 | ||
| 512 | 64.49 | 57.11 | 67.02 | 66.25 | 51.80 | 65.15 | 53.32 | 61.29 | 68.24 | ||
| GPT-2 | Acc | 32 | 70.53 | 75.99 | 69.32 | 70.00 | 51.02 | 71.69 | 68.59 | 71.88 | 75.42 |
| 64 | 74.41 | 75.76 | 73.77 | 74.04 | 54.32 | 73.20 | 71.12 | 71.88 | 78.92 | ||
| 128 | 79.77 | 75.77 | 80.18 | 80.93 | 59.66 | 79.44 | 71.74 | 71.88 | 82.58 | ||
| 512 | 84.07 | 75.86 | 84.76 | 84.89 | 71.59 | 84.30 | 71.74 | 74.06 | 85.75 | ||
| F1 | 32 | 66.57 | 72.45 | 64.89 | 69.89 | 48.45 | 71.19 | 65.50 | 70.00 | 75.10 | |
| 64 | 73.91 | 70.87 | 72.32 | 73.94 | 53.87 | 69.79 | 66.58 | 70.97 | 78.88 | ||
| 128 | 79.49 | 71.16 | 80.00 | 80.79 | 59.48 | 76.10 | 66.13 | 71.88 | 82.54 | ||
| 512 | 84.01 | 75.56 | 84.72 | 84.86 | 70.42 | 83.88 | 66.13 | 74.64 | 85.72 | ||
| GPT-3.5 | Acc | 32 | 90.54 | 92.55 | 92.44 | 92.85 | 61.82 | 93.27 | 84.42 | 89.10 | 95.80 |
| 64 | 96.85 | 91.00 | 96.86 | 97.32 | 77.70 | 95.76 | 82.58 | 89.65 | 98.01 | ||
| 128 | 97.50 | 91.60 | 98.00 | 98.00 | 92.54 | 96.26 | 85.33 | 89.85 | 98.06 | ||
| 512 | 98.97 | 92.60 | 98.99 | 98.92 | 98.13 | 98.05 | 85.57 | 90.62 | 99.14 | ||
| F1 | 32 | 90.27 | 92.71 | 92.42 | 92.81 | 60.95 | 92.72 | 84.43 | 89.76 | 95.80 | |
| 64 | 96.84 | 91.49 | 96.86 | 97.47 | 77.33 | 95.45 | 86.16 | 89.92 | 98.01 | ||
| 128 | 97.50 | 91.96 | 98.00 | 98.00 | 92.50 | 97.57 | 86.13 | 89.77 | 98.06 | ||
| 512 | 98.85 | 92.71 | 98.93 | 98.92 | 98.13 | 97.88 | 86.20 | 90.62 | 99.14 | ||
| HC3 | Acc | 32 | 93.36 | 97.30 | 95.33 | 95.46 | 77.00 | 92.11 | 94.54 | 87.70 | 97.19 |
| 64 | 96.97 | 98.13 | 97.81 | 97.81 | 91.69 | 95.50 | 95.03 | 88.87 | 98.59 | ||
| 128 | 97.56 | 98.29 | 98.17 | 98.14 | 95.43 | 97.57 | 95.10 | 88.87 | 98.63 | ||
| 512 | 98.85 | 98.31 | 98.93 | 98.99 | 97.98 | 98.58 | 95.13 | 90.62 | 99.15 | ||
| F1 | 32 | 93.34 | 97.30 | 95.32 | 95.45 | 76.47 | 92.07 | 94.29 | 88.39 | 97.19 | |
| 64 | 96.97 | 98.12 | 97.81 | 97.81 | 91.67 | 95.50 | 94.95 | 89.92 | 98.59 | ||
| 128 | 97.56 | 98.29 | 98.17 | 98.14 | 95.43 | 97.59 | 95.01 | 89.92 | 98.63 | ||
| 512 | 98.85 | 98.31 | 98.93 | 98.99 | 97.98 | 98.59 | 95.05 | 91.06 | 99.15 |